OpenStudy (dimples97):

Can someone please help me with this question? Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (2, 16)

OpenStudy (lapinblanc):

The coefficient in front of the equation's x is your slope. So what you do is use the following equation:\[y-y _{1}=m(x-x _{1})\] where m is your slope. So, use the slope you know, the points you were given input them into the above equation Solve for y Voila!

OpenStudy (trojanpoem):

The line is parallel the equation is given in the form y= mx + b m = 1 (slope) parallel m1 = m2 m2 = m1 = 1 y = x + b 16 = 2 + b b = 14 y = x + 14

OpenStudy (lapinblanc):

Parallel lines have the exact same slope.